OpenStudy (anonymous):

Please Help Find the probability of rolling a 12 with a pair of dice.

OpenStudy (anonymous):

Give two fair six-sided die there are 36 possible combinations. How many of those combinations equal 12?

OpenStudy (anonymous):

Would it be 12/36?

OpenStudy (anonymous):

Listing the possible combinations can be tedious. But, it is the best way to make sure you don't miss something. If you started listing possible combinations, how many of those combinations equals 12? A table is very helpful for problems like this. First Die 2nd Die Sum 1 1 2 1 2 3 1 3 4 etc...

OpenStudy (anonymous):

The only way to get a 12 is to get two sixes. The chance of rolling a six is 1/6 and chance of rolling two sixes is \[\frac{ 1}{ 6}*\frac{ 1 }{ 6 }=\frac{ 1 }{ 36 }\]